# Early Stages of Self-Healing at Tungsten Grain Boundaries from Ab Initio Machine Learning Simulations

### v1

@@ Abstract
Nanostructured tungsten has been reported as a possible alternative plasma-facing material due to its potential ability to self-heal radiation-induced defects, a property that is attributed to its high density of grain boundaries (GB). Here, we study the initial stages of self-healing at tungsten interfaces with molecular dynamics simulations driven by a machine-learning interatomic potential tailored to one of the most common GBs found in experiments. Our model accurately reproduces the ab initio potential energy surface derived from density functional theory (DFT) calculations and outperforms previously reported empirical interatomic potentials in predicting defect energetics. The simulations reveal low-temperature defect migration to GBs driven by rapid dumbbell-like ordering and subsequent accommodation along GB grooves. In contrast to empirical potentials, which predict unexpected GB degradation at high temperatures after defect migration, our model maintains stable GB motifs over the investigated temperature range. The temperature-dependent defect counts, evaluated using an Arrhenius-like fit, yield an average interstitial migration energy of 0.048 eV, in agreement with experiment.
@@ 1 Introduction
Irradiation of coarse-grained materials by particles (e.g., ions, neutrons) with energies exceeding the displacement threshold, as occurs in the inertial nuclear reactors operated in the direct drive approach, produces atomic displacements that generate vacancies and self-interstitial atoms (SIAs). At a given temperature, both these types of defects migrate towards defect sinks such as grain boundaries (GBs), dislocations and voids. However, in general, they do so at quite different energies and time scales. These differences mean that they usually do not meet and recombine, leading to a permanent degradation of material properties such as electrical conductivity and mechanical response [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. From an atomistic point of view, the situation can be different if vacancies and SIAs are forced to be very close to each other, which can be done by reducing the grain size (nanostructuring) or, in other words, by introducing a high density of GBs. In this case, the recombination probability between vacancies and SIAs increases, leading to their mutual annihilation at the surface or GBs, a phenomenon commonly referred to as self-healing, so that the material spontaneously returns to its unirradiated structure. Nevertheless, this self-healing behavior only occurs under specific conditions, well captured phenomenologically by the three-temperature (low, medium, and high) model proposed by Beyerlein et al. [13]. In the low temperature regime, vacancies are immobile, so any damage inflicted on the nanostructured material is greater than to its coarse-grained counterpart [14, 15]. This is because highly mobile SIAs migrate along the bulk toward the GBs, leaving immobile vacancies in the grain interior. At intermediate temperatures, where vacancies are still immobile, some of the radiation-induced damage can be annihilated by the re-emission of SIAs from the GBs [15, 16, 17]. Finally, at high temperatures, vacancies become mobile and move toward GBs, where they effectively recombine with SIAs, annihilating the radiation-induced damage. The evolution of defects and the temperature threshold for moving from one temperature regime to another depend on the material composition and microstructure.
@@
Even though this model describes the origin of the self-healing behavior, its effectiveness depends strongly on the GB architecture [18]. Therefore, a proper description of the self-healing capabilities of nanostructured materials requires an understanding of how defects migrate and interact with the GB at the atomic scale, which can be achieved by performing computer simulations. Traditionally, two main approaches have been used: ab initio molecular dynamics (AIMD) — typically based on Density Functional Theory (DFT) to compute interatomic forces — and classical Molecular Dynamics (MD) using empirical or semi-empirical interatomic potentials (hereafter IPs). While AIMD offers high precision in capturing atomic interactions, it is constrained by limited time and length scales [19, 20]. Conversely, classical MD allows the simulation of larger systems over longer timescales, but relies on IPs that often have limited accuracy and transferability [21, 22]. As a result, an IP developed for coarse-grained materials may not be sufficiently accurate to adequately describe radiation-induced damage in nanostructured materials. Indeed, this fact could limit self-healing studies, where the simulation of defect migration requires an extremely precise description of GB energetics as well as relatively large simulation boxes and long simulation times. Recent years have witnessed the rapid development of machine learning interatomic potentials (MLIPs) that leverage large datasets, often derived from quantum-mechanical calculations, to build flexible, highly accurate models of atomic interactions [23, 24, 25]. These advanced methods have already been applied to challenging problems such as radiation damage cascades, surface reactions, and phase transformations under extreme thermomechanical conditions [26, 27, 28, 29]. Critically, they can achieve near-ab initio accuracy at a fraction of the computational cost, thus enabling simulations on time and length scales relevant to fusion reactor environments.
@@
In this work, we present a computational study of how intrinsic defects migrate and interact with a tungsten GB, with an eye towards having a deeper understanding of the self-healing in nanostructured tungsten, one of the candidates proposed as a plasma-facing material in nuclear fusion reactors [30, 31, 32, 33, 34]. For this purpose, we develop a new IP, dubbed MLIP-W, based on the Deep Potential methodology [35] and trained on quantum mechanical DFT data. Although some MLIPs for tungsten have been developed in recent years [36, 37, 38], none of them has been specifically trained to describe GBs and thus are not guaranteed to preserve ab initio accuracy for complex environments involving both defects and interfaces. We apply our model to study the early stages of self-healing wherein SIAs migrate to GBs. Note that vacancy migration occurs on a timescale of seconds, far beyond the timescales accessible in MD simulations, and for this reason it is not considered in our study. We explicitly train our potential to accurately describe the commonly observed experimental W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ GB over a wide temperature range, as well as the interaction of this GB with point defects. We then compare the performance of MLIP-W with widely used empirical potentials and highlight its advantages. Moreover, we perform automated defect identification and classification using a fingerprinting toolkit that considers the local atomic environments [39, 40]. Our results show that the MLIP-W outperforms conventional empirical IPs in capturing defect migration, and it also preserves the structure of the GB at high temperatures, in contrast with empirical IPs. These findings highlight the effectiveness of the MLIP-W in describing self-healing near GBs, offering a more accurate computational framework for modeling radiation damage.
@@ 2.1 Development of the MLIP-W
To accurately capture the ab initio potential energy surface (PES) for W, we used the smooth version of the Deep Potential method [35, 41] as implemented in DeePMD-kit [42, 43], an open-source deep learning software for training neural network potentials. In our workflow, we coupled this methodology with an active learning scheme to iteratively refine the training set and generate an extensive set of atomic configurations and their corresponding DFT-derived energies and forces [44, 45]. This methodology has been successfully demonstrated on several occasions, yielding excellent results [46, 47]. At each step of the active learning process, short MD simulations were run to identify "high-error" configurations based on force discrepancies among a set of independently trained networks. Then, we recalculated these configurations using DFT, and retrained the model. For a more detailed explanation of the active learning approach, see Subsection 4.1. In our case, the dataset includes configurations for bulk W and for the W110delimited-⟨⟩110$\langle$ 110$\rangle$ 110 ⟩/W112delimited-⟨⟩112$\langle$ 112$\rangle$ 112 ⟩ GB, each thermalized at temperatures ranging from 300 K to 1000 K to cover the full range of interest and ensure robust performance at high temperatures. In addition, vacancies and SIAs were randomly introduced to capture a variety of environments beyond the pristine bulk and the dynamics of these defects, including recombination and migration to the GB, were simulated for 100 ps. The diversity of these configurations in the training dataset is visualized in Fig. 1 A, which shows a t-distributed stochastic neighbor embedding (t-SNE) representation of a subset of 10,000 descriptor vectors (DVs)–numerical representations of local atomic environments–drawn from the full set. Distinct clusters are observed: light purple accounts for bulk-like environments, dark purple for SIA defects, beige for atoms near free surfaces or vacancies, and pale pink for a variety of defect-rich regions, including the GB.
Figure 1: Diversity and accuracy of the MLIP-W training dataset. (A) t-SNE representation of local atomic environments in the final training set using DeePMD-derived descriptors. The four clusters correspond to near-bulk atoms (light purple), SIA-type defects, e.g. $\langle 111\rangle$ dumbbells (dark purple), free surfaces or vacancies (beige), and defect environments, including GBs (pale pink). (B) and (C) Parity plots comparing ML-predicted total energy and forces with DFT values. Insets show error distributions ($\epsilon_{E}$, $\epsilon_{F}$) with near-Gaussian fits for energies and Lorentzian-like shapes for forces.
@@
To measure its overall accuracy, we computed the root mean square (RMS) errors for energies and forces over the entire training set, obtaining $\epsilon_{E}^{\mathrm{RMS}}\approx 0.84$ meV/atom and $\epsilon_{F}^{\mathrm{RMS}}\approx 151$ meV/Å. These errors compare favorably with other state-of-the-art MLIPs [48, 49]. Figs. 1 B and C show parity plots of the ML energies and forces against the DFT references, with insets showing their respective error distributions. In both cases, the distributions are centered near zero. This performance highlights the ability of the model to capture subtle features of the PES, especially in systems where the geometry deviates significantly from the bulk, such as defects, GBs, and surfaces. Note that the reference DFT calculations were performed using the PBE exchange-correlation functional. Therefore, the accuracy of the derived potential remains inherently limited by this specific choice.
@@ 2.2 Structural properties and defect energetics benchmarks
Table 1 shows a comparison of several fundamental properties of BCC W as predicted by the MLIP-W, DFT (computed in this work), experimental measurements (where available), and two widely used empirical IPs, namely, an Embedded Atom Method (EAM) (Marinica et al.[50]) and a Tersoff (Juslin et al.[51]). We find that the MLIP-W closely reproduces the DFT-calculated lattice parameter and vacancy formation energy, which are also in good agreement with the experimental results. In contrast, the EAM potential slightly underestimates the lattice parameter, while the Tersoff potential yields a larger vacancy formation energy. Furthermore, both the MLIP-W and EAM potentials accurately reproduce the formation energies for SIAs as determined by the reference DFT calculations, whereas the Tersoff potential consistently predicts significantly lower formation energies relative to DFT. A similar trend is observed for di-vacancies, i.e., the MLIP-W predicts negative binding energies for second-neighbor (2nn) di-vacancies, in agreement with DFT, whereas both the Tersoff and EAM potentials yield positive values for that quantity. These results illustrate that the MLIP-W is able to accurately capture defect energetics, which is critical for subsequent simulations focused on defect evolution. However, while these static benchmarks provide an initial validation, and even parameterizations for larger time and length scale models, they alone do not necessarily guarantee accurate defect kinetics or correct long-term microstructural evolution.
@@
Table 1: Tungsten lattice parameter ($a_{0}$), formation energy for a single vacancy ($E_{f}$) and various interstitial configurations including dumbbells in the $\langle 110\rangle$ ($E_{f,db110}$) and $\langle 111\rangle$ ($E_{f,db111}$) directions, and in an octahedral site ($E_{f,octa}$), and di-vacancy binding energies in first nearest neighbors ($E_{b}$(V-V) 1nn) and second nearest neighbors ($E_{b}$(V-V) 2nn) arrangements. We compare results obtained experimentally and calculated using DFT, MLIP-W (this work), Tersoff [51], and EAM potentials [50].
Property Exp. DFT MLIP-W Tersoff [51] EAM [50]
$a_{0}$ (BCC) (Å) 3.17 [52, 53, 54] 3.17 3.16 3.16 3.14
$E_{f}$(V) (eV) 3.15 [55] 3.34 3.30 3.71 3.49
$E_{f,db110}$ (eV) 11.04 10.97 10.16 10.86
$E_{f,db111}$ (eV) 10.76 10.86 9.50 10.40
$E_{f,octa}$ (eV) 13.11 12.80 10.39 12.73
$E_{b}$(V-V) 1nn (eV) 0.7 [56] 0.41 0.29 0.47 0.49
$E_{b}$(V-V) 2nn (eV) -0.19 -0.21 0.41 0.38
Another parameter to be considered is the Generalized Stacking Fault Energy (GSFE) which is a key material property that characterizes the energy profile associated with dislocation motion in a crystal lattice. It is well known that in BCC tungsten, the {110} plane is a primary slip plane with the unstable stacking fault energy reflecting the energy barrier to dislocation motion. Meanwhile, analyzing the GSFE along the {112} plane provides evidence that twinning is less likely to dominate the deformation process in W, indicating that dislocation slip remains the primary mode of plastic deformation. The GSFE profiles for the {110} and {112} planes obtained using the MLIP-W and compared to traditional empirical IPs [50, 51, 57] and DFT [58] results are shown in the Supplementary Figs. S1 B and C, respectively. Although the MLIP-W was not explicitly trained on GSFE predictions, it shows good agreement with DFT data in terms of the barrier height, even though it does not fully capture the asymmetric shape of the GSFE profile for the {110} plane. Fig. S2 D shows the mapping of the GSFE across the fault plane.

### v2

@@ Abstract
Nanostructured tungsten has been reported as a possible alternative plasma-facing material due to its potential ability to self-heal radiation-induced defects, a property that is attributed to its high density of grain boundaries (GB). Here, we study the initial stages of self-healing at tungsten interfaces with molecular dynamics simulations driven by a machine-learning interatomic potential tailored to one of the most common GBs found in experiments. Our model accurately reproduces the ab initio potential energy surface derived from density functional theory (DFT) calculations and outperforms previously reported empirical and machine learning interatomic potentials in predicting defect energetics. The simulations reveal low-temperature defect migration to GBs driven by rapid dumbbell-like ordering and subsequent accommodation along GB grooves. In contrast to empirical potentials, which predict unexpected GB degradation at high temperatures after defect migration, our model maintains stable GB motifs over the investigated temperature range. The temperature-dependent defect counts, evaluated using an Arrhenius-like fit, yield an average interstitial migration energy of 0.048 eV, in agreement with experiment.
@@ 1 Introduction
Irradiation of coarse-grained materials by particles (e.g., ions, neutrons) with energies exceeding the displacement threshold, as occurs in the inertial nuclear reactors operated in the direct drive approach, produces atomic displacements that generate vacancies and self-interstitial atoms (SIAs). At a given temperature, both types of defects migrate towards defect sinks such as grain boundaries (GBs), dislocations, and voids. However, in general, they do so at quite different energies and time scales. These differences mean that they usually do not meet and recombine, leading to a permanent degradation of material properties such as electrical conductivity and mechanical response [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. From an atomistic point of view, the situation can be different if vacancies and SIAs are forced to be very close to each other, which can be done by reducing the grain size (nanostructuring) or, in other words, by introducing a high density of GBs. In this case, the recombination probability between vacancies and SIAs increases, leading, under certain conditions, to their mutual annihilation at the surface or GBs, a phenomenon commonly referred to as self-healing, so that the material spontaneously returns to its unirradiated structure. Nevertheless, as previously mentioned, this self-healing behavior only occurs under specific conditions, well captured phenomenologically by the three-temperature (low, medium, and high) model proposed by Beyerlein et al. [13]. In the low temperature regime, vacancies are immobile, so any damage inflicted on the nanostructured material is greater than on its coarse-grained counterpart [14, 15]. This is because highly mobile SIAs migrate along the bulk toward the GBs, leaving immobile vacancies in the grain interior. At intermediate temperatures, where vacancies are still immobile, some of the radiation-induced damage can be annihilated by the re-emission of SIAs from the GBs [15, 16, 17]. Finally, at high temperatures, vacancies become mobile and move toward GBs, where they effectively recombine with SIAs, annihilating the radiation-induced damage. The evolution of defects and the temperature threshold for moving from one temperature regime to another depend on the material composition and microstructure.
Even though this model describes the origin of the self-healing behavior, its effectiveness depends strongly on the GB architecture [18]. Therefore, a proper description of the self-healing capabilities of nanostructured materials requires an understanding of how defects migrate and interact with the specific GB at the atomic scale, which can be achieved by performing computer simulations.
@@
Traditionally, two main approaches have been used: ab initio molecular dynamics (AIMD) — typically based on Density Functional Theory (DFT) to compute interatomic forces — and classical Molecular Dynamics (MD) using empirical or semi-empirical interatomic potentials (hereafter IPs). While AIMD offers high precision in capturing atomic interactions, it is constrained by limited time and length scales [19, 20]. Conversely, classical MD allows the simulation of larger systems over longer timescales, but relies on IPs that often have limited accuracy and transferability [21, 22]. As a result, an IP developed for coarse-grained materials may not be sufficiently accurate to adequately describe radiation-induced damage in nanostructured materials. Indeed, this fact could limit self-healing studies, where the simulation of defect migration requires an extremely precise description of GB energetics as well as relatively large simulation boxes and long simulation times.
Recent years have witnessed the rapid development of machine learning interatomic potentials (MLIPs) that leverage large datasets, often derived from quantum-mechanical calculations, to build flexible, highly accurate models of atomic interactions [23, 24, 25]. These advanced methods have already been applied to challenging problems such as radiation damage cascades, surface reactions, and phase transformations under extreme thermomechanical conditions [26, 27, 28, 29]. Critically, they can achieve near-ab initio accuracy at a fraction of the computational cost, thus enabling simulations on time and length scales relevant to capturing the early stages of self-healing at tungsten interfaces.
@@
In this work, we present a computational study of how intrinsic defects migrate and interact with a tungsten GB, with an eye towards having a deeper understanding of the self-healing in nanostructured tungsten, one of the candidates proposed as a plasma-facing material in nuclear fusion reactors [30, 31, 32, 33, 34]. We focus on atomistic damage since the thermomechanical effects that PFMs will undergo and that could lead to significant temperature increases and, consequently, grain growth (if this is not somehow stabilized), could be alleviated by using a/or by combining different strategies: lower target yield, employing radiation mitigation strategies, and optimizing the chamber geometry. However, atomistic effects cannot be avoided. To study the latter, we develop a IP, dubbed MLIP-W, based on the Deep Potential methodology [35] and trained on quantum mechanical DFT data. Although some MLIPs for tungsten have been developed in recent years [36, 37, 38], none of them have been specifically trained to describe GBs meaning that they may not be guaranteed to preserve ab initio accuracy for complex environments involving both defects and interfaces. We apply our model to study the early stages of self-healing wherein SIAs migrate to GBs. Because the nanosecond-scale MD simulations carried out here are orders of magnitude shorter than the second-scale vacancy diffusion times, vacancy motion cannot be resolved and is not considered in our study. We explicitly train our potential to accurately describe the $\langle 110\rangle/W$$\langle 112\rangle$ GB over a wide temperature range, as well as the interaction of this GB with point defects. It is worth mentioning that we select this particular GB because it is the most frequently observed in sputter-deposited tungsten films, a fabrication route that can be easily scaled up for industrial production [39]. We then compare the performance of MLIP-W to two recent, state-of-the-art MLIPs, as well as to widely used empirical potentials, to highlight its advantages. Moreover, we perform automated defect identification and classification using a fingerprinting toolkit that considers the local atomic environments [40, 41]. Our results show that the MLIP-W outperforms existing IPs in capturing defect migration. Moreover, the MLIP-W also preserves the structure of the GB at high temperatures, in contrast with previous IPs. These findings show the effectiveness of the MLIP-W in describing self-healing near GBs, offering a more accurate computational framework for modeling radiation damage.
@@ 2.1 Development of the MLIP-W
To accurately capture the ab initio potential energy surface (PES) for W, we used the smooth version of the Deep Potential method [35, 42] as implemented in DeePMD-kit [43, 44], an open-source deep learning software for training neural network potentials. In our workflow, we coupled this methodology with an active learning scheme to iteratively refine the training set and generate an extensive set of atomic configurations and their corresponding DFT-derived energies and forces [45, 46]. This methodology has been successfully demonstrated on several occasions, yielding excellent results [47, 48]. At each step of the active learning process, short MD simulations were run to identify "high-error" configurations based on force discrepancies among a set of independently trained networks. These uncertain configurations were recalculated using DFT, and the model was retrained accordingly. A detailed explanation of the active learning approach is provided in Subsection 4.1.
In our case, the initial training dataset includes bulk W configurations and the $\langle 110\rangle/W$$\langle 112\rangle$ GB, with random atomic displacements applied to sample regions of the PES near equilibrium positions (see the left panels of Fig. 1). To accurately capture intrinsic defect behavior, we randomly introduced vacancies and SIAs into both the bulk and GB configurations. However, to further explore GB-defect interactions at finite temperatures, we followed an active learning approach to identify and include new atomic environments critical to our study. The central panel of Fig. 1 presents a t-distributed stochastic neighbor embedding (t-SNE) visualization of these environments. This visualization was generated from a subset of 10,000 descriptor vectors (DVs)—which are numerical representations of local atomic environments—drawn from the full training set. The environments identified and integrated via active learning are shown in gray. They typically correspond to highly complex scenarios involving multiple defect interactions, such as interstitial atoms interacting with GBs or free surfaces, as shown by the dashed circles in the panel on the right. The integration of active learning with the flexible MLIP framework substantially improves upon the limitations of empirical potentials, which are characterized by a closed-form function, enabling the treatment of highly nontrivial defect-GB scenarios.
Figure 1: Development of the MLIP-W. The left panels depict representative atomic environments included in the initial training dataset, which is composed of configurations generated by random displacements and single point defects. At the top of the center panel, a simplified flowchart outlines the active-learning loop used to enrich the dataset. Below that, a t-SNE visualization of local atomic environments encoded by DeePMD descriptors reveals the following: bulk-like environments (beige), interstitial defects (purple), free surface atoms (orange), GB atoms (red), and environments sampled via active learning (gray). The panel on the right provides an example of the complexity and diversity of the atomic environments in the dataset by showing the presence of multiple defect types in a single configuration. The error distributions and parity plots at the bottom quantify the overall accuracy achieved after approximately 50 active learning cycles.
To ensure the comprehensiveness of our dataset and to minimize potential biases or gaps, we continued the active learning process iteratively, completing a total of 50 cycles, until no new configurations exhibiting uncertainties above our predefined threshold were identified in subsequent MD simulations. Moreover, convergence in the descriptor distribution was observed, with new configurations consistently falling inside the existing descriptor space. This indicates that no new significant atomic environments were discovered in the later cycles of active learning.
@@
To measure its overall accuracy, we computed the root mean square (RMS) errors for energies and forces over the entire training set, obtaining $\epsilon_{E}^{\mathrm{RMS}}\approx 0.84$ meV/atom and $\epsilon_{F}^{\mathrm{RMS}}\approx 151$ meV/Å. These errors compare favorably to other state-of-the-art MLIPs [49, 50]. The bottom panels of Fig. 1 include the error distributions for energies and forces on the left, followed by parity plots comparing the ML- and DFT-predicted energies and forces, respectively, on the right. This performance highlights the ability of the model to capture subtle features of the PES, especially in systems where the geometry deviates significantly from the bulk, such as defects, GBs, and surfaces. Note that the reference DFT calculations were performed using the PBE exchange-correlation functional. Therefore, the accuracy of the derived potential remains inherently limited by this specific choice.
@@ 2.2 Structural properties and defect energetics benchmarks
Table  1 compares several fundamental properties of BCC W as predicted by our MLIP-W model, DFT (computed in this work), experiments (where available), two widely used empirical IPs—an embedded atom method (EAM) potential by Marinica et al. [51] and a Tersoff potential by Juslin & Wirth [52], and by two recent MLIPs—the tabulated Gaussian Approximation Potential (TabGAP) by Byggmästar et al. [38] and the Spectral Neighbor Analysis Potential (SNAP) by Wood et al. [53]. Our model closely matches DFT results for all properties and outperforms other models. MLIP-W predicts the single vacancy formation energy, $E_{f}(\mathrm{V})$, with a deviation of $0.04$ eV with regards to the DFT values, and the $\langle 110\rangle$ and $\langle 111\rangle$ dumbbell formation energies, $E_{f,db110}$ and $E_{f,db111}$, with deviations of $0.07$ eV and $0.10$ eV, respectively. The octahedral interstitial formation energy, $E_{f,octa}$, is underestimated by $0.31$ eV relative to DFT. Although the last deviation is larger than the others, the value calculated with the MLIP-W remains the closest to the DFT reference. The EAM potential shows the next smallest discrepancy, of 0.38 eV. For the first-nearest-neighbor di-vacancy binding energy, $E_{b}(V\!-\!V\ 1\mathrm{nn})$, the Tersoff (+0.06+0.06 eV), EAM (+0.08+0.08 eV) and SNAP (+0.07+0.07 eV) potentials produce values closer to the DFT reference than MLIP-W (0.12-0.12 eV). However, MLIP-W is the only model, apart from SNAP, which gives a value of 0.07-0.07 eV, to reproduce the negative second-nearest-neighbor di-vacancy binding energy, $E_{b}(V\!-\!V\ 2\mathrm{nn})$, reporting 0.21-0.21 eV versus 0.19-0.19 eV in DFT. We also evaluated the generalized stacking fault energy (GSFE) profiles, which are critical for dislocation-mediated defect motion. Though MLIP-W was not trained on GSFE data, we found that it closely reproduces DFT barrier heights on both the {110} and {112} planes (see Supplementary Figures S1B–C) [51, 52, 54, 55]. This agreement confirms the reliability of MLIP-W regarding the energetics that govern defect migration. For full details, see Supplementary Note 1.
@@
Table 1: Tungsten lattice parameter ($a_{0}$), formation energy for a single vacancy ($E_{f}$) and various interstitial configurations including dumbbells in the $\langle 110\rangle$ ($E_{f,db110}$) and $\langle 111\rangle$ ($E_{f,db111}$) directions, and in an octahedral site ($E_{f,octa}$), and di-vacancy binding energies in first nearest neighbors ($E_{b}$(V–V) 1nn) and second nearest neighbors ($E_{b}$(V–V) 2nn) arrangements. We compare results obtained experimentally and calculated using DFT, MLIP-W (this work), Tersoff [52], EAM [51], TabGAP [38], and SNAP [53].
Property Exp. DFT MLIP-W Tersoff EAM TabGAP SNAP
$a_{0}$ (BCC) (Å) 3.17 [56, 57, 58] 3.17 3.16 3.16 3.14 3.18 3.18
$E_{f}$(V) (eV) 3.15 [59] $3.34$ 3.30 3.71 3.49 3.22 3.20
$E_{f,db110}$ (eV) $11.04$ 10.97 10.16 10.86 10.64 9.76
$E_{f,db111}$ (eV) $10.76$ 10.86 9.50 10.40 10.34 9.55
$E_{f,octa}$ (eV) $13.11$ 12.80 10.39 12.73 10.50 11.5
$E_{b}$(V–V) 1nn (eV) 0.7 [60] $0.41$ 0.29 0.47 0.49 0.06 0.48
$E_{b}$(V–V) 2nn (eV) 0.19-0.19 0.21-0.21 0.41 0.38 0.05 0.07-0.07
However, while these static benchmarks provide an initial validation, and even parameterizations for larger time and length scale models, they alone do not necessarily guarantee accurate defect kinetics or correct long-term microstructural evolution. The next essential step is to evaluate the performance of the MLIP-W and the other potentials under finite-temperature conditions that are directly relevant to our simulations of defect-GB interactions and their dynamics. To this end, we took the $\langle 110\rangle/W$$\langle 112\rangle$ GB described above, randomly introduced interstitials and vacancies, and evolved the system using our model at 300, 650, and 1000 K to obtain representative configurations. We computed atomic forces from selected snapshots of these dynamic simulations using MLIP-W, as well as empirical (EAM and Tersoff) and ML-based (TabGAP and SNAP) potentials. We then compared these forces with DFT reference forces (see Fig. 2). This approach allows us to rigorously test the predictive accuracy of the potentials under realistic simulation conditions and ensure reliable modeling of defect kinetics at finite temperatures.
Figure 2: Finite-temperature force accuracy of the evaluated potentials. Parity plots comparing the atomic forces predicted by each IP (vertical axis) with the reference DFT forces (horizontal axis). The plots are for $\langle 110\rangle/W$$\langle 112\rangle$ GB configurations containing Frenkel pairs. The configurations were sampled after MD evolution at 300 K (top row), 650 K (middle row), and 1000 K (bottom row). The columns correspond to MLIP-W, Tersoff, EAM, TabGAP, and SNAP from left to right. The solid black line shows perfect correlation, and each panel reports the RMSE.
Fig.  2 confirms that MLIP-W provides the most accurate forces at all the three investigated temperatures. Its RMSE increases only modestly from 0.114 eV Å-1 at 300 K to 0.202 eV Å-1 at 1000 K. However, it remains lower than that of any other model at every temperature. The next-best performer, TabGAP, exhibits RMSE values between 0.332 and 0.411 eV Å-1, which is roughly a factor of two higher than MLIP-W in the worst case scenario. SNAP and Tersoff give comparable, larger errors ($\approx 0.40\textendash 0.64$ eV Å-1), with Tersoff consistently the less accurate of the two. EAM performs the worst: its RMSE increases from 0.863 eV Å-1 at 300 K to 1.137 eV Å-1 at 1000 K, more than five times the MLIP-W error at that temperature. These trends demonstrate that MLIP-W achieves nearly DFT-level force accuracy, particularly in capturing GB-defect interactions, even at the high temperatures explored in this study.